OpenStudy (anonymous):

in base 13 6 * 9 = ?

OpenStudy (april115):

isn't that 54?

OpenStudy (anonymous):

it is 54 in base 10 not in base 13 try again

OpenStudy (april115):

Ohh :O

OpenStudy (april115):

42

OpenStudy (anonymous):

54 = (4*13) + 2 = 2 (mod 13)

OpenStudy (anonymous):

good job you 2

OpenStudy (april115):

Because in base 13 6 times 9 it happens to equal 42 <--- its just a Coincidence. http://www.6times9.com